$5900 s invested, part of it at 12% and part of it at 8%. For a certain year, the total yield is $596.00. How? much was invested at each rate
Answer by ptfile(81)

Let x = the amount invested at 12%
y = the amount invested at 8%
$5900 is invested.
x+y=5900
part of it at 12% and part of it at 8%. For a certain year, the total yield is $596.
.12x+.08y=596
.12x(-12.5)+.08y(-12.5)=596(-12.5)
-1.5x- y=-7450
x+ y= 5900
-1.5x- y=-7450
-.5x=-1550
x= 3100
Substitute x with $3100
$3100+y=5900
y=5900-3100
y=2800
Out of $5900, $3100 was invested at 12% and $2800 was invested at 8%.
Check
.12(3100)+.08(2800)=596
372 + 224 =596
596 =596